# Efficient Test Data Generation for MC/DC with OCL and Search

### v1

@@ 4.1.3 Identifying Conflicting MC/DC Constraints
Since a conflicting MC/DC constraint cannot be solved, our strategy identifies the conflict by finding the dependent clauses first. The dependency among clauses is identified by examining the use of common primitive attributes. If two or more clauses are dependent, we analyze the fitness values for half of the total search budget. For a non-conflicting constraint, its fitness values should improve with the increase in fitness evaluations regardless of constraint complexity. In the case where fitness values do not improve for a specified search budget, the constraint is considered conflicting and the search process is terminated. For the confirmation, we need to manually analyze whether the constraint was conflicting or not. Besides conflict due to dependent clauses, a constraint with independent clauses can be conflicting. In our approach, such conflict can be identified by analyzing the fitness values. However, tools like QMaxUSE [108], or techniques such as [27] and  [67] can be used to detect such conflicts.
After the reformulation of all OCL constraints for the MC/DC criterion, the next step is to solve MC/DC constraints. A straightforward approach is to solve each MC/DC constraint individually (e.g., using [4]). The problem with this approach is that it is a time-consuming task. It can be observed from the reformulation step that when a single OCL constraint is reformulated to apply MC/DC, one constraint is converted to multiple MC/DC constraints (as shown in Listing 2). According to our experience with industrial constraints [56], it is very time-consuming to solve industrial-scale OCL constraints with different complexity levels [5]. It can take up to 15 minutes to solve a single OCL constraint. Thus, solving each MC/DC constraint separately can consume a lot of computing resources (such as time and search budget [54]), especially, when the number of clauses in an OCL constraint increases. To handle this problem, we propose two methods to improve the search for solving MC/DC constraints. In the following sections, we discuss both methods.
@@ 4.2 Solving MC/DC Constraints using Case-based Reasoning
Analyzing the example MC/DC constraints as shown in Listing 2, we notice that constraints C1, C2, and C3 have the same first clause, and constraints C3 and C4 have an identical third clause. Whenever MC/DC is applied, the resulting MC/DC constraints exhibit this behavior. So, rather than solving each of the MC/DC constraints individually, we can reuse a component of already solved constraints. Therefore, when the previous solution is reused, it contains the solution for an already solved clause(s). As a result, the search algorithm starts the search from a partially solved individual rather than a random individual. In this way, finding a solution for a constraint becomes much easier as compared to solving a constraint from scratch.
For this purpose, we adopt the concept of case-based reasoning (CBR) [64] to solve MC/DC constraints. The main reason is to reuse the previous solution while solving the current MC/DC constraint. The CBR approach provides a well-suited method to reuse solutions of previously solved problems while solving a current problem that has a part similar to that of the previously solved problems.
@@
$d\left(c_{t_{j}},\ c_{s_{i}j}\right)=\left\{\begin{array}[]{l}1,\ \ c_{t_{j}}=\ c_{s_{i}j}\\ 0,\ \ c_{t_{j}}\neq\ c_{s_{i}j}\end{array}\right.$ (2)
@@
Using the Equations 1 and 2, the similarity score of each solution in the repository with the target predicate is calculated. Based on the similarity score, we get the possible set of closest solutions.
@@
$S_{T}\mathrm{=}\left\{\begin{array}[]{l}S_{c}\mathrm{,\ \ }f_{s_{c}}\mathrm{<}f_{s_{r}}\\ S_{r}\mathrm{,\ \ }f_{s_{c}}\geq f_{s_{r}}\end{array}\right.$ (3)
@@ 4.3 Solving MC/DC Constraints using Range Reduction
We address the problem of using a fixed range for all attributes. We present heuristics to select the range for a specific constraint with Integer and Real type of attributes. In the case of Boolean and Enumeration type of attributes, the range is already small. The Boolean attribute can have only two values in its search domain i.e., true and false. The Enumeration type of attribute contains a fixed number of Enumeration literals. For example, the enumeration for weekdays contains only seven Enumeration literals. In addition, constraint clauses may contain OCL collection operations and user-defined operations. However, such operations are rarely used in industrial constraints [56, 5]. Moreover, the search space in such operations is typically small and can be solved easily without range reduction. Therefore, we focus on two commonly used primitive types i.e., Integer and Real. In the following, we discuss heuristics to select the range based on the nature of OCL constraints.
An OCL constraint consists of several clauses. A clause contains an OCL expression. In the case of Integer and Real type attributes, a clause is represented by an expression with an attribute name, a relational operator, and a constant. For example, OCL constraint C2 as shown in Listing 3 contains an invariant with a single clause i.e., ‘self.mission.waypoints $>$ self.mission.MIN_WP_LIMIT’. In this clause, the ‘waypoints’ is the attribute name, the relational operator is ‘$>$’, and the constant is MIN_WP_LIMIT. In the following, we present heuristics to reduce the range for different types of clauses with Integer and Real type attributes.
@@ 4.3.1 Constraint with Single Clause
$A_{min}\mathrm{=}\left\{\begin{array}[]{l}(c-c)*s_{f}\mathrm{,\ \ }c\mathrm{>}0\\ (c+c)*s_{f}\mathrm{,\ \ }c\mathrm{<}0\\ -1*s_{f}\mathrm{,\ \ }c\mathrm{=}0\end{array}\right.$ (4)
$A_{max}\mathrm{=}\left\{\begin{array}[]{l}(c+c)*s_{f}\mathrm{,\ \ }c\mathrm{>}0\\ (c-c)*s_{f}\mathrm{,\ \ }c\mathrm{<}0\\ 1*s_{f}\mathrm{,\ \ }c\mathrm{=}0\end{array}\right.$ (5)
In Equations 4 and 5, $s_{f}$ is a scaling factor. Its values are positive non-zero integers i.e., $s_{f}=\{1,2,3,\dots\}$. The values of $s_{f}$ determine the scale of minimum and maximum values for specifying the range. Domain experts need to select scaling factor value considering constraints complexity. The value of the scaling factor needs be to set at the start of the search. The higher the scaling factor, the higher will be the range.
For example, consider a constraint C2 as shown in Listing 1. The constraint C2 contains an invariant with a single clause i.e., ‘self.mission.waypoints $>$ self.mission.MIN_WP_LIMIT’. In this clause, waypoints is the attribute (A) and the constant c is MIN_WP_LIMIT. If the value of MIN_WP_LIMIT is 10, the constant c is 10, i.e., c=10. If the value of the scaling factor is set to 1 i.e., $s_{f}=1$, the minimum and maximum values are calculated as $A_{min}=10-10*1=0$ and $A_{max}=10+10*1=20$. Therefore, the range of the attribute (A) with $s_{f}=1$ is $[0,10]$. If the value of the scaling factor is set to 2 i.e., $s_{f}=2$, the minimum and maximum values are calculated as $A_{min}=10-10*2=-10$ and $A_{max}=10+10*2=30$. Therefore, the range of the attribute (A) with $s_{f}=2$ is $[-10,30]$. As the value of the scaling factor increases, the range of the attribute increases.
@@
For example, consider a constraint C3 as shown in Listing 1. The constraint C3 contains an invariant with a single clause i.e., ‘self.mission.waypoints $>$ self.mission.MIN_WP_LIMIT+100’. In this clause, waypoints is the attribute (A) and there are two constants i.e., $c_{1}=$MIN_WP_LIMIT and $c_{2}=100$. The constants expression contains an addition arithmetic operator (i.e., +). If the value of constant MIN_WP_LIMIT is 20, the resultant constant after solving the arithmetic expression is $c=120$. If the value of the scaling factor is set to 1 i.e., $s_{f}=1$, the minimum and maximum values are calculated as $A_{min}=120-120*1=0$ and $A_{max}=120+120*1=240$. Therefore, the range of the attribute (A) with $s_{f}=1$ is $[0,240]$. If the value of the scaling factor is set to 2 i.e., $s_{f}=2$, the minimum and maximum values are calculated as $A_{min}=120-120*2=-120$ and $A_{max}=120+120*2=360$. Therefore, the range of the attribute (A) with $s_{f}=2$ is $[-120,360]$.
@@
For the example constraint C4 as shown in Listing 1, $A_{1}$ and $A_{2}$ represent the attribute distance and the variable minDist respectively. If the smallest positive integer value selected at random is 5, i.e., $c=5$. If the value of the scaling factor is set to 1 i.e., $s_{f}=1$, the minimum and maximum values are calculated as $A_{1_{min}}=A_{2_{min}}=5-5*1=0$ and $A_{1_{max}}=A_{2_{max}}=5+5*1=10$. Therefore, the range of $A_{1}$ and $A_{2}$ with $s_{f}=1$ is $[0,10]$. If the value of the scaling factor is set to 2 i.e., $s_{f}=2$, the minimum and maximum values are calculated as $A_{1_{min}}=A_{2_{min}}=5-5*2=-5$ and $A_{1_{max}}=A_{2_{max}}=5+5*2=15$. Therefore, the range of $A_{1}$ and $A_{2}$ with $s_{f}=2$ is $[-5,15]$.
@@
For example, consider a constraint C5 as shown in Listing 1. The constraint contains a pre-condition with a single clause i.e., ‘self.distance+1000 $>$ minDist’. In this clause, distance is the attribute (A), the constant c is 1000, and minDist is the variable. If the value of the scaling factor is set to 1 i.e., $s_{f}=1$, the minimum and maximum values are calculated as $A_{min}=1000-1000*1=0$ and $A_{max}=1000+1000*1=2000$. Therefore, the range of the attribute (A) with $s_{f}=1$ is $[0,2000]$. If the value of the scaling factor is set to 2 i.e., $s_{f}=2$, the minimum and maximum values are calculated as $A_{min}=1000-1000*2=-1000$ and $A_{max}=1000+1000*2=3000$. Therefore, the range of the attribute (A) with $s_{f}=2$ is $[-1000,3000]$.
Case 4: Multiple attributes/variables on both sides of a relational operator. OCL constraints can contain some clauses in which multiple attributes are used on both sides of a relational operator. If there is no constant on either side of a relational operator, choose the smallest positive integer value randomly in the range [0, 100]. If there is one constant involved in a clause, the value of the constant c is used. If there are more than one constants involved in a clause, the minimum value from constants is used as a resultant constant c. After determining the constant value, the range is calculated using Equations 4 and 5.
For example, consider a constraint C6 as shown in Listing 1. The constraint contains a pre-condition with a single clause i.e., ‘self.distance+1500 $>$ minDist+maxDist’. In this clause, distance is the attribute, the constant c is 1500, and minDistnt and maxDist are the variables. If the value of the scaling factor is set to 1 i.e., $s_{f}=1$, the minimum and maximum values are calculated as $A_{min}=1500-1500*1=0$ and $A_{max}=1500+1500*1=3000$. Therefore, the range of the attribute (A) with $s_{f}=1$ is $[0,3000]$. If the value of the scaling factor is set to 2 i.e., $s_{f}=2$, the minimum and maximum values are calculated as $A_{min}=1500-1500*2=-1500$ and $A_{max}=1500+1500*2=4500$. Therefore, the range of the attribute (A) with $s_{f}=2$ is $[-1500,4500]$.

### v2

@@ 4.1.3 Identifying Conflicting MC/DC Constraints
Since a conflicting MC/DC constraint cannot be solved, our strategy identifies the conflict by finding the dependent clauses first. The dependency among clauses is identified by examining the use of common primitive attributes. If two or more clauses are dependent, we analyze the fitness values for half of the total search budget. For a non-conflicting constraint, its fitness values should improve with the increase in fitness evaluations regardless of constraint complexity. In the case where fitness values do not improve for a specified search budget, the constraint is considered conflicting and the search process is terminated. For the confirmation, we need to manually analyze whether the constraint was conflicting or not. Besides conflict due to dependent clauses, a constraint with independent clauses can be conflicting. In our approach, such conflict can be identified by analyzing the fitness values. However, tools like QMaxUSE [108], or other techniques [27, 67] can be used to detect such conflicts.
After the reformulation of all OCL constraints for the MC/DC criterion, the next step is to solve MC/DC constraints. A straightforward approach is to solve each MC/DC constraint individually (e.g., using the approach by Ali et al. [4]). The problem with this approach is that it is a time-consuming task. It can be observed from the reformulation step that when a single OCL constraint is reformulated to apply MC/DC, one constraint is converted to multiple MC/DC constraints (as shown in Listing 2). According to our experience with industrial constraints [56], it is very time-consuming to solve industrial-scale OCL constraints with different complexity levels [5]. It can take up to 15 minutes to solve a single OCL constraint. Thus, solving each MC/DC constraint separately can consume a lot of computing resources (such as time and search budget [54]), especially, when the number of clauses in an OCL constraint increases. To handle this problem, we propose two methods to improve the search for solving MC/DC constraints. In the following sections, we discuss both methods.
@@ 4.2 Solving MC/DC Constraints using Case-based Reasoning
Analyzing the example MC/DC constraints as shown in Listing 2, we notice that constraints C1, C2, and C3 have the same first clause, and constraints C3 and C4 have an identical third clause. Whenever MC/DC is applied, the resulting MC/DC constraints exhibit this behavior. So, rather than solving each MC/DC constraint individually, we can reuse a component of already solved constraints. Therefore, when the previous solution is reused, it contains the solution for an already solved clause(s). As a result, the search algorithm starts the search from a partially solved individual rather than a random individual. In this way, finding a solution for a constraint becomes much easier compared to solving a constraint from scratch.
For this purpose, we adopt the concept of case-based reasoning (CBR) [64] to reuse the previous solution while solving MC/DC constraints. The CBR approach provides a well-suited method to reuse solutions of previously solved problems while solving a current problem with a part similar to the previously solved problems.
@@
$d\left(c_{t_{j}},\ c_{s_{i}j}\right)=\left\{\begin{array}[]{ l l }1&\quad c_{t_{j}}=\ c_{s_{i}j}\\ 0&\quad c_{t_{j}}\neq\ c_{s_{i}j}\end{array}\right.$ (2)
@@
Using Equations 1 and 2, the similarity score of each solution in the repository with the target predicate is calculated. Based on the similarity score, we get the possible set of closest solutions.
@@
$S_{T}\mathrm{=}\left\{\begin{array}[]{ l l }S_{c}&\quad f_{s_{c}}\mathrm{<}f_{s_{r}}\\ S_{r}&\quad f_{s_{c}}\geq f_{s_{r}}\end{array}\right.$ (3)
@@ 4.3 Solving MC/DC Constraints using Range Reduction
We address the problem of using a fixed range for all attributes. We present heuristics to select the range for a specific constraint with Integer and Real type of attributes. In the case of Boolean and Enumeration type of attributes, the range is already small. The Boolean attribute can have only two values in its search domain, i.e., true and false. The Enumeration type of attribute contains a fixed number of Enumeration literals. For example, the enumeration for weekdays contains only seven Enumeration literals. In addition, constraint clauses may contain OCL collection operations and user-defined operations. However, such operations are rarely used in industrial constraints [56, 5]. Moreover, the search space in such operations is typically small and can be solved easily without range reduction. Therefore, we focus on two commonly used primitive types, i.e., Integer and Real. In the following, we discuss heuristics to select the range based on the nature of OCL constraints.
An OCL constraint consists of several clauses. A clause contains an OCL expression. In the case of Integer and Real type attributes, a clause is represented by an expression with an attribute name, a relational operator, and a constant. For example, OCL constraint C2 as shown in Listing 3 contains an invariant with a single clause, i.e., ‘self.mission.waypoints $>$ self.mission.MIN_WP_LIMIT’. In this clause, the ‘waypoints’ is the attribute name, the relational operator is ‘$>$’, and the constant is MIN_WP_LIMIT. In the following, we present heuristics to reduce the range for different types of clauses with Integer and Real type attributes. First, we introduce cases involving single clause constraints (Section 4.3.1), followed by a discussion of the cases involving multiple clauses joined with Boolean operators (Section 4.3.2). At the single clause level, cases in Section 4.3.1 can be considered special cases of those presented in Section 4.3.2. We describe each case individually to ensure a more in-depth understanding of every scenario.
@@ 4.3.1 Constraint with Single Clause
$A_{min}\mathrm{=}\left\{\begin{array}[]{ l l }{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0}&\quad c\mathrm{>}0\\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2*c}*s_{f}&\quad c\mathrm{<}0\\ -1*s_{f}&\quad c\mathrm{=}0\end{array}\right.$ (4)
$A_{max}\mathrm{=}\left\{\begin{array}[]{ l l }{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2*c}*s_{f}&\quad c\mathrm{>}0\\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0}&\quad c\mathrm{<}0\\ 1*s_{f}&\quad c\mathrm{=}0\end{array}\right.$ (5)
In Equations 4 and 5, $s_{f}$ is a scaling factor. Its values are positive non-zero integers, i.e., $s_{f}=\{1,2,3,\dots\}$. The values of $s_{f}$ determine the scale of minimum and maximum values for specifying the range. Initially, domain experts must select a scaling factor value considering the constraints’ complexity and case study specifications, which is a one-time effort. This predetermined scaling factor value should be set at the beginning of the search process. It can then be utilized for subsequent generation tasks. It is important to note that a higher scaling factor will result in a correspondingly higher range.
For example, consider a constraint C2 as shown in Listing 1. The constraint C2 contains an invariant with a single clause, i.e., ‘self.mission.waypoints $>$ self.mission.MIN_WP_LIMIT’. In this clause, waypoints is the attribute (A) and the constant c is MIN_WP_LIMIT. If the value of MIN_WP_LIMIT is 10, the constant c is 10, i.e., c=10. If the value of the scaling factor is set to 1, i.e., $s_{f}=1$, the minimum and maximum values are calculated as $A_{min}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0}$ and $A_{max}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2*10*1=20}$. Therefore, the range of the attribute (A) with $s_{f}=1$ is $[0,20]$. If the value of the scaling factor is set to 2, i.e., $s_{f}=2$, the minimum and maximum values are calculated as $A_{min}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0}$ and $A_{max}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2*10*2=40}$. Therefore, the range of the attribute (A) with $s_{f}=2$ is $[0,40]$. As the value of the scaling factor increases, the range of the attribute increases.
@@
For example, consider a constraint C3 as shown in Listing 1. The constraint C3 contains an invariant with a single clause, i.e., ‘self.mission.waypoints $>$ self.mission.MIN_WP_LIMIT+100’. In this clause, waypoints is the attribute (A) and there are two constants i.e., $c_{1}=$MIN_WP_LIMIT and $c_{2}=100$. The constants expression contains an addition arithmetic operator (i.e., +). If the value of constant MIN_WP_LIMIT is 20, the resultant constant after solving the arithmetic expression is $c=120$. If the value of the scaling factor is set to 1, i.e., $s_{f}=1$, the minimum and maximum values are calculated as $A_{min}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0}$ and $A_{max}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2*120*1=240}$. Therefore, the range of the attribute (A) with $s_{f}=1$ is $[0,240]$. If the value of the scaling factor is set to 2, i.e., $s_{f}=2$, the minimum and maximum values are calculated as $A_{min}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0}$ and $A_{max}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2*120*2=480}$. Therefore, the range of the attribute (A) with $s_{f}=2$ is $[0,480]$.
@@
For the example constraint C4 as shown in Listing 1, $A_{1}$ and $A_{2}$ represent the attribute distance and the variable minDist respectively. If the smallest positive integer value selected at random is 5, i.e., $c=5$. If the value of the scaling factor is set to 1, i.e., $s_{f}=1$, the minimum and maximum values are calculated as $A_{1_{min}}=A_{2_{min}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0}$ and $A_{1_{max}}=A_{2_{max}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2*5*1=10}$. Therefore, the range of $A_{1}$ and $A_{2}$ with $s_{f}=1$ is $[0,10]$. If the value of the scaling factor is set to 2, i.e., $s_{f}=2$, the minimum and maximum values are calculated as $A_{1_{min}}=A_{2_{min}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0}$ and $A_{1_{max}}=A_{2_{max}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2*5*2=20}$. Therefore, the range of $A_{1}$ and $A_{2}$ with $s_{f}=2$ is $[0,20]$.
@@
For example, consider a constraint C5 as shown in Listing 1. The constraint contains a pre-condition with a single clause, i.e., ‘self.distance+1000 $>$ minDist’. In this clause, distance is the attribute (A), the constant c is 1000, and minDist is the variable. If the value of the scaling factor is set to 1, i.e., $s_{f}=1$, the minimum and maximum values are calculated as $A_{min}=A_{2_{max}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0}$ and $A_{max}=A_{2_{max}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2*1000*1=2000}$. Therefore, the range of the attribute (A) with $s_{f}=1$ is $[0,2000]$. If the value of the scaling factor is set to 2, i.e., $s_{f}=2$, the minimum and maximum values are calculated as $A_{min}=A_{2_{max}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0}$ and $A_{max}=A_{2_{max}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2*1000*2=4000}$. Therefore, the range of the attribute (A) with $s_{f}=2$ is $[0,4000]$.
Case 4: Multiple attributes/variables on both sides of a relational operator. OCL constraints can contain clauses in which multiple attributes are used on both sides of a relational operator. If there is no constant on either side of a relational operator, choose the smallest positive integer value randomly in the range [0, 100]. If there is one constant involved in a clause, the value of the constant c is used. If there are more than one constants involved in a clause, the minimum value from constants is used as a resultant constant c. After determining the constant value, the range is calculated using Equations 4 and 5.
For example, consider a constraint C6 as shown in Listing 1. The constraint contains a pre-condition with a single clause, i.e., ‘self.distance+1500 $>$ minDist+maxDist’. In this clause, distance is the attribute, the constant c is 1500, and minDist and maxDist are the variables. If the value of the scaling factor is set to 1, i.e., $s_{f}=1$, the minimum and maximum values are calculated as $A_{min}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0}$ and $A_{max}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2*1500*1=3000}$. Therefore, the range of the attribute (A) with $s_{f}=1$ is $[0,3000]$. If the value of the scaling factor is set to 2, i.e., $s_{f}=2$, the minimum and maximum values are calculated as $A_{min}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0}$ and $A_{max}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2*1500*2=6000}$. Therefore, the range of the attribute (A) with $s_{f}=2$ is $[0,6000]$.